# Searching for leptophilic composite asymmetric dark sector at $e^{+}e^{-}$ colliders

### v1

@@ Abstract
Composite asymmetric dark matter (ADM) models provide a well-motivated paradigm that simultaneously explains dark matter (DM) relic density and matter-antimatter asymmetry. In these models, the mass of the DM candidate (the lightest dark baryon) is generated through the dark confinement scale dynamics. Although the leptophilic composite ADM model offers a viable framework, comprehensive studies of its collider phenomenology are absent. This work systematically explores novel signatures from leptophilic composite asymmetric dark sector at both low-energy and high-energy $e^{+}e^{-}$ colliders as well as other existing collider constraints. We demonstrate detectability of TeV-scale mediators along with sub-GeV to GeV-scale lightest dark mesons at Belle II and its proposed far detector, GAZELLE, as well as CEPC experiments. Moreover, these experiments exhibit complementary coverage of the model parameter space.
@@ I Introduction
The particle nature of dark matter (DM) and the origin of matter-antimatter asymmetry remain outstanding challenges in particle physics, astrophysics, and cosmology. The proposed DM models span mass scales from ultralight to tens of solar masses Battaglieri:2017aum ; Lin:2019uvt ; Ferreira:2020fam ; Belenchia:2021rfb , with well-motivated candidates including: weakly interacting massive particles (WIMPs) Arcadi:2017kky ; Arcadi:2024ukq , strongly interacting massive particles (SIMPs) Hochberg:2014dra , hidden sector DM Battaglieri:2017aum , WIMPless DM Feng:2008ya , asymmetric DM (ADM) Petraki:2013wwa ; Zurek:2013wia , freeze-in DM Bernal:2017kxu , QCD axions Baryakhtar:2025jwh , and primordial black holes Villanueva-Domingo:2021spv . Despite extensive searches, no conclusive evidence supports any candidate to date. Consequently, diverse experimental strategies have been developed to probe their properties.
@@
Composite ADM Blinnikov:1982eh ; Blinnikov:1983gh ; Khlopov:1989fj ; Alves:2009nf ; SpierMoreiraAlves:2010err ; Bai:2013xga ; Ibe:2018juk ; Ibe:2018tex ; Ibe:2019ena ; Hall:2019rld ; Zhang:2021orr ; Bottaro:2021aal ; Ibe:2021gil ; Ritter:2022opo represents a prominent scenario in which the dark sector features a hidden $SU(N)$ gauge symmetry analogous to SM QCD. In such models, accidental baryon number conservation stabilizes the lightest dark baryon, making it a natural DM candidate. Crucially, the DM mass is directly related to the dark confinement scale $\Lambda_{d}$, thereby being dynamically determined rather than a free parameter. Mediators facilitate dark sector and SM sector coupling to generate shared particle-antiparticle asymmetry. We adopt the framework in Refs. Bai:2013xga ; Zhang:2021orr where bi-charged scalar mediator decays produce asymmetries in both sectors.
@@
Two bi-charged scalar mediators exist in the literature: baryophilic Bai:2013xga and leptophilic Zhang:2021orr types, both capable of simultaneously explaining DM relic density and matter-antimatter asymmetry. The baryophilic mediator carries charges under both hidden $SU(N)$ and SM QCD symmetries, coupling to a dark quark and an SM quark. This enables copious dark quark production at the LHC where, given the dark quark mass is much smaller than the collision energy, the produced dark quarks become energetic and highly boosted. When the scale of $\Lambda_{d}$ is also much smaller than the collision energy, subsequent QCD-like showering and hadronization generate distinctive dark jets that facilitate novel search strategies at the large hadron collider (LHC) Schwaller:2015gea ; Cohen:2015toa ; Cohen:2017pzm ; Park:2017rfb ; Renner:2018fhh ; Mies:2020mzw ; Linthorne:2021oiz ; Archer-Smith:2021ntx ; Carrasco:2023loy ; Carmona:2024tkg . In contrast, the leptophilic mediator is coupled with a dark quark and an SM lepton, suppressing LHC production rates as noted in Zhang:2021orr . Therefore, lepton colliders consequently provide ideal environments for probing such models. Although Ref. Zhang:2021orr proposed preliminary studies, comprehensive analyzes remain lacking. This work therefore systematically explores leptophilic composite ADM signatures on low-energy and high-energy $e^{+}e^{-}$ colliders.
@@
Collider searches for leptophilic scalar mediators resemble searches of slepton in supersymmetry models when dark hadrons inside dark jets escape detection at LHC and LEP ATLAS:2019lff ; CMS:2020bfa ; CMS:2022syk ; CMS:2024gyw ; ATLAS:2024fub ; Fox:2011fx . We therefore rescale existing slepton constraints to these mediators. While collision energies exceeding the dark confinement scale $\Lambda_{d}$ can produce dark showers, the intensity of showering depends critically on energy: higher energies generate denser parton showers. At low-energy $e^{+}e^{-}$ colliders such as $B$-factories, sparse dark showers enable tracking of individual displaced dark pions and kinematic reconstruction. Similar approaches have been studied for the dark photon mediator Bernreuther:2022jlj , though signatures differ fundamentally from this work. Conversely, at high-energy $e^{+}e^{-}$ colliders (CEPC, FCC-ee, ILC), dense dark showers necessitate identification of displaced, collimated lepton jets rather than individual dark pions Zhang:2021orr . When dark jets become undetectable, mono-photon signatures Liu:2019ogn become relevant. This work establishes complementary coverage of parameter space through these distinct search strategies.
@@
This paper is structured as follows: Section II reviews simplified leptophilic composite ADM models. Section III examines existing collider constraints. Novel dark shower signatures at $B$-factories and CEPC are explored in Sections IV and V, respectively. Section VI presents our conclusions and future research directions.
@@ II A brief review of QCD-like dark sector with bi-charged mediators
In the leptophilic composite ADM models proposed in Ref. Zhang:2021orr , new heavy particles, including Majorana fermions and $SU(3)_{D}$ charged Dirac fermions, generate out-of-equilibrium processes and explicitly break hidden baryon number conservation. However, these particles lie significantly above current collider energy scales and are consequently decoupled from our analysis. We therefore restrict the Lagrangian to the relevant scalar mediator $X$ and dark quark $q_{d}$ interactions as follows [^1] The possible Higgs portal coupling between the SM Higgs doublet field and the $X$ field is irrelevant to this study. Therefore, we assume that its contributions are negligible and can be neglected here.:
@@
Figure 1: Diagrams illustrating dark shower (DS) production evolving into invisible or emerging lepton jets. Left panel: $e^{+}e^{-}$ annihilation via $t$-channel mediator $X$ producing a dark quark pair ($q_{d}\overline{q_{d}}$). Right panel: $q\overline{q}$ annihilation yielding mediator pairs, where each mediator decays to a dark quark and a SM lepton.
At lepton and hadron colliders, dark quarks are produced through $t$-channel processes and/or mediator decays, as shown in Fig. 1. Post-production, they acquire significant kinetic energy ($E\gg\Lambda_{d}$), where $\Lambda_{d}$ denotes the dark confinement scale. Assuming the dark QCD sector mirrors SM QCD dynamics, these dark quarks undergo dark parton shower evolution and hadronization, forming a collinear bunch of dark mesons.
In subsequent sections, we adopt the parameter settings for the QCD-like dark sector from Refs. Schwaller:2015gea ; Mies:2020mzw ; Archer-Smith:2021ntx ; Carrasco:2023loy to enable concrete simulation predictions. We assume three dark colors ($N_{C_{d}}=3$) and seven dark flavors ($N_{f_{d}}=7$), with the following mass relations: $\Lambda_{d}=m_{q_{d}}=2m_{\pi_{d}}$, and $m_{\rho_{d}}=4m_{\pi_{d}}$. Note that the dark confinement scale $\Lambda_{d}$ is closely related to the mass of the lightest dark baryon, which serves as the DM candidate Bai:2013xga ; Schwaller:2015gea ; Renner:2018fhh ; Mies:2020mzw ; Zhang:2021orr . All dark quarks are taken to be mass-degenerate, while dark pions and dark rho mesons are similarly degenerate in both flavor diagonal and off-diagonal states. Other dark baryons and mesons, assumed substantially heavier than dark pions and rho mesons, exhibit negligible production rates from dark showers and hadronization due to mass suppression. We therefore exclude them from this analysis.
Integrating out the heavy mediator (final terms in Eq. (1)) yields a dimension-six operator,
@@
which enables dark pion ($\pi_{d}$) decays to charged lepton pairs. The decay width is Schwaller:2015gea :
$\Gamma(\pi_{d}\to\bar{\ell}\ell)=\frac{\kappa^{4}f_{\pi_{d}}^{2}m_{\ell}^{2}m_{\pi_{d}}}{32\pi m_{X}^{4}},\quad\text{with}\quad\ell=e,\mu,\tau.$ (3)
Here $f_{\pi_{d}}$ is the dark pion decay constant and $m_{\ell}$ the lepton mass. Parameterizing the proper lifetime as
$c\tau_{0}=\frac{c\hbar}{\Gamma_{\pi_{d}}}\approx 1.92\,\text{m}\times\left(\frac{1}{\kappa^{4}}\right)\left(\frac{1\,\text{GeV}}{f_{\pi_{d}}}\right)^{2}\left(\frac{0.1\,\text{GeV}}{m_{\ell}}\right)^{2}\left(\frac{1\,\text{GeV}}{m_{\pi_{d}}}\right)\left(\frac{m_{X}}{1\,\text{TeV}}\right)^{4},$ (4)
reveals that dark pions propagate measurable distances from the primary vertex (PV) before decaying into charged lepton pairs.
@@
To illustrate the typical proper decay length of $\pi_{d}$ versus $m_{\pi_{d}}$, we fix $\kappa=1.0$, $f_{\pi_{d}}=m_{\pi_{d}}$, and $m_{X}=1$ TeV as benchmark values. In order to simplify our discussions here, the lepton flavor university (LFU) is assumed such that the common coupling $\kappa$ applies to the interactions of mediator with a dark quark and a SM charged lepton for each lepton flavor. Figure 2 shows that for $2m_{e}<m_{\pi_{d}}<2m_{\mu}$, $\pi_{d}$ decays to electron-positron pairs. The proper decay length exceeds $10^{7}$ m, allowing electrons and positrons to escape detection. When $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$, the muon decay channel opens, with $\pi_{d}$ predominantly decaying into muon pairs. The resulting decay length $\sim 1$ m makes these long-lived $\pi_{d}$ detectable. Finally, for $m_{\pi_{d}}>2m_{\tau}$, the tau channel becomes accessible, and $\pi_{d}$ primarily decays to tau pairs, potentially exhibiting prompt decays.
@@ III Existing collider constraints on model parameters
Figure 2 reveals a dramatic change in the proper decay length of $\pi_{d}$ at the $m_{\pi_{d}}\geq 2m_{\mu}$ threshold. We therefore classify two dark pion mass ranges: $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$ and $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, and discuss their respective existing collider constraints.
@@
Figure 3: Exclusion region for dark pion masses $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$. The red dashed line separates long-lived (below) and prompt (above) $X^{\pm}$ decays. The blue and black dashed lines indicate where all $\pi_{d}$ particles escape detection at the LHC and LEP, respectively. The blue and black shaded regions correspond to slepton search constraints (LHC) ATLAS:2019lff and mono-photon search constraints (LEP) Fox:2011fx ; Liu:2019ogn . The green region represents the parameter space where the proper lifetime of $\pi_{d}$ exceeds 1 second, conservatively evaluated at $m_{\pi_{d}}=0.2\text{ GeV}$.
@@
Figure 4: Exclusion region for dark pion masses $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$. The blue dashed line indicates where all $\pi_{d}$ particles escape detection at the LHC, making this region subject to slepton search constraints. The red dashed line retains the same interpretation as in Fig. 3. The green region represents the parameter space where the proper lifetime of $\pi_{d}$ exceeds 1 second, conservatively evaluated at $m_{\pi_{d}}=3\text{ GeV}$.
@@
First, we consider the production of scalar mediator pairs through $pp\to X^{+}X^{-}$, with decays $X^{+}\to\ell^{+}q_{d}$ and $X^{-}\to\ell^{-}\overline{q_{d}}$ ($\ell=e,\mu,\tau$). Here, $\kappa$ in Eq. (1) affects only the $X^{\pm}$ lifetime, not the production cross-section. When $X^{\pm}$ decays promptly and dark parton showers are invisible at ATLAS and CMS detectors, signatures resemble slepton and stau searches in supersymmetry models ATLAS:2019lff ; ATLAS:2024fub . Under the LFU assumption, we recast these constraints onto this model. We generate a UFO model file for Eq. (1) via FeynRules Alloul:2013bka and simulate signal processes with MadGraph5_aMC@NLO Alwall:2014hca (version 3.5.8). For each $m_{X}$, we scan $\kappa$ and define the mediator prompt decay region by $\gamma\beta c\tau_{X}<0.1\text{ mm}$ in the lab frame. Figures 3 and 4 show regions above the red dashed line ($\kappa\gtrsim\mathcal{O}(10^{-7}-10^{-6})$) for the mediator prompt decays recast from slepton and stau constraints [^2] The signal production cross-sections rescaled to match Fig. 3 of Ref. Zhang:2021orr with QCD NLO+NLL corrections at $13\text{ TeV}$ LHC Fiaschi:2018xdm .. Additionally, because there are plenty of dark pions produced after dark showers and hadronization, we require the $\pi_{d}$ proper decay length (without any boost) larger than $10\text{ m}$ to conservatively assume all dark pions escape ATLAS and CMS detectors as missing energy. Specifically, we set $m_{\pi_{d}}=f_{\pi_{d}}=0.2\text{ GeV}$ in Fig. 3 and $m_{\pi_{d}}=f_{\pi_{d}}=3\text{ GeV}$ in Fig. 4 to estimate the upper bounds of $\kappa$. Regions below blue dashed lines ($\kappa\lesssim\mathcal{O}(1-10)$ for $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$; $\kappa\lesssim\mathcal{O}(0.1-1)$ for $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$) correspond to invisible dark showers which are adapted for recasting.
@@
Under the LFU assumption, the mediator $X^{\pm}$ has approximately equal decay branching ratios of $1/3$ to each flavor channel ($e^{\pm}q_{d}$, $\mu^{\pm}q_{d}$, $\tau^{\pm}q_{d}$). Consequently, we include an extra factor of $1/9$ when applying stau pair production constraints. Comparison with ATLAS results from Table 01 of auxiliary material in Ref. ATLAS:2024fub shows that for $m_{X}>200$ GeV and the dark quark mass $m_{q_{d}}=1$ GeV, the effective cross-section is at least a factor of two below the experimental bounds. Thus, no exclusion region exists for $m_{X}>200$ GeV, though stau constraints exclude $m_{X}\lesssim 160$ GeV.
@@
For slepton pair production constraints in Fig. 3c of Ref. ATLAS:2019lff , the analysis involves all combinations of a pair of leptons from electrons and muons in the final states. Therefore, we apply a factor of $4/9$ in our process. Notably, while the constraints include contributions from both $\tilde{\ell}^{+}_{L}\tilde{\ell}^{-}_{L}$ and $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ production, our signal process specifically corresponds to the $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ channel. We therefore isolate the right-handed slepton constraints using the fraction of $pp\to\tilde{\ell}^{+}_{L}\tilde{\ell}^{-}_{L}$ and $pp\to\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ production cross-sections from Refs. Fuks:2013lya ; Fiaschi:2018xdm . The resulting effective cross-section in our process exceeds ATLAS exclusion limits for $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ in the mass range $m_{X}\in[200,700]$ GeV. Additionally, after accounting for the average Lorentz boost of the mediator ($\gamma\beta\simeq 1.5$), we estimate the prompt decay boundaries within this mass range.
@@
Conversely, for $\pi_{d}$ with moderate decay lengths, the novel signatures of displaced lepton jets are expected to be observable at the LHC. However, Ref. Zhang:2021orr notes detection challenges for the $pp\to X^{+}X^{-}$ production channel due to: (1) insufficient production cross-sections, and (2) detection efficiencies after considering the whole cut-flow below $\sim 0.1$. Moreover, significant irreducible cosmic-ray backgrounds further complicate the observation. While LHC Run 3 and HL-LHC may improve sensitivity, current ATLAS displaced lepton jet searches ATLAS:2022izj exclusively target Higgs exotic decays, distinct from this model’s signature of two opposite-sign leptons plus two displaced lepton jets. Consequently, recasting strategies from Ref. ATLAS:2022izj are inapplicable. A dedicated analysis (beyond the scope of this work) would be required, though we establish that the LHC constraints remain weaker than those derived for $e^{+}e^{-}$ colliders in Sections IV and V.

### v2

@@ Abstract
Composite asymmetric dark matter (ADM) models provide a well-motivated paradigm that simultaneously explains dark matter (DM) relic density and matter-antimatter asymmetry. In these models, the mass of the DM candidate (the lightest dark baryon) is generated through the dark confinement scale dynamics. Although the leptophilic composite ADM model offers a viable framework, comprehensive studies of its collider phenomenology are absent. This work systematically explores novel signatures from the leptophilic composite asymmetric dark sector at both low-energy and high-energy $e^{+}e^{-}$ colliders as well as other existing collider constraints. We demonstrate detectability of TeV-scale mediators along with sub-GeV to GeV-scale lightest dark mesons at Belle II and its proposed far detector, GAZELLE, as well as Circular Electron Positron Collider/Future Circular Collider-electron-positron experiments. Moreover, these experiments exhibit complementary coverage of the model parameter space.
@@ I Introduction
The particle nature of dark matter (DM) and the origin of matter-antimatter asymmetry remain outstanding challenges in particle physics, astrophysics, and cosmology. The proposed DM models span mass scales from ultralight to tens of solar masses Battaglieri:2017aum ; Lin:2019uvt ; Ferreira:2020fam ; Belenchia:2021rfb , with well-motivated candidates including weakly interacting massive particles (WIMPs) Arcadi:2017kky ; Arcadi:2024ukq , strongly interacting massive particles Hochberg:2014dra , hidden sector DM Battaglieri:2017aum , WIMPless DM Feng:2008ya , asymmetric DM (ADM) Petraki:2013wwa ; Zurek:2013wia , freeze-in DM Bernal:2017kxu , QCD axions Baryakhtar:2025jwh , and primordial black holes Villanueva-Domingo:2021spv . Despite extensive searches, no conclusive evidence supports any candidate to date. Consequently, diverse experimental strategies have been developed to probe their properties.
@@
Composite ADM Blinnikov:1982eh ; Blinnikov:1983gh ; Khlopov:1989fj ; Alves:2009nf ; SpierMoreiraAlves:2010err ; Bai:2013xga ; Ibe:2018juk ; Ibe:2018tex ; Ibe:2019ena ; Hall:2019rld ; Zhang:2021orr ; Bottaro:2021aal ; Ibe:2021gil ; Ritter:2022opo represents a prominent scenario in which the dark sector features a hidden $SU(N)$ gauge symmetry analogous to SM QCD. In such models, accidental dark baryon number conservation stabilizes the lightest dark baryon, making it a natural DM candidate. Crucially, the DM mass is directly related to the dark confinement scale $\Lambda_{d}$, thereby being dynamically determined rather than a free parameter. Mediators facilitate dark sector and SM sector coupling to generate shared particle-antiparticle asymmetry. We adopt the framework in Refs. Bai:2013xga ; Zhang:2021orr , where bi-charged scalar mediator decays produce asymmetries in both sectors.
@@
Two bi-charged scalar mediators exist in the literature: baryophilic Bai:2013xga and leptophilic Zhang:2021orr types, both capable of simultaneously explaining DM relic density and matter-antimatter asymmetry. The baryophilic mediator carries charges under both hidden $SU(N)$ and SM QCD symmetries, coupling to a dark quark and an SM quark. This enables copious dark quark production at the LHC where, given the dark quark mass is much smaller than the collision energy, the produced dark quarks become energetic and highly boosted. When the scale of $\Lambda_{d}$ is also much smaller than the collision energy, subsequent QCD-like showering and hadronization generate distinctive dark jets that facilitate novel search strategies at the LHC Schwaller:2015gea ; Cohen:2015toa ; Cohen:2017pzm ; Park:2017rfb ; Renner:2018fhh ; Mies:2020mzw ; Linthorne:2021oiz ; Archer-Smith:2021ntx ; Carrasco:2023loy ; Carmona:2024tkg ; Liu:2025bbc . In contrast, the leptophilic mediator is coupled with a dark quark and an SM lepton, suppressing LHC production rates as noted in Zhang:2021orr . Therefore, lepton colliders consequently provide ideal environments for probing such models. Although Ref. Zhang:2021orr proposed preliminary studies, comprehensive analyses remain lacking. This work, therefore, systematically explores leptophilic composite ADM signatures on low-energy and high-energy $e^{+}e^{-}$ colliders.
@@
Collider searches for leptophilic scalar mediators resemble searches of slepton in supersymmetry models when dark hadrons inside dark jets escape detection at the LHC and LEP ATLAS:2019lff ; CMS:2020bfa ; CMS:2022syk ; CMS:2024gyw ; ATLAS:2024fub ; Fox:2011fx . We, therefore, rescale existing slepton constraints to these mediators. While collision energies exceeding the dark confinement scale $\Lambda_{d}$ can produce dark showers, the intensity of showering depends critically on energy: higher energies generate denser parton showers. At low-energy $e^{+}e^{-}$ colliders such as $B$ factories, sparse dark showers enable tracking of individual displaced dark pions and kinematic reconstruction. Similar approaches have been studied for the dark photon mediator Bernreuther:2022jlj , though signatures differ fundamentally from this work. Conversely, at high-energy $e^{+}e^{-}$ colliders [Circular Electron Positron Collider(CEPC), Future Circular Collider-electron-positron(FCC-ee), and ILC], dense dark showers necessitate identification of displaced, collimated lepton jets rather than individual dark pions Zhang:2021orr . When dark jets become undetectable, monophoton signatures Liu:2019ogn become relevant. This work establishes complementary coverage of parameter space through these distinct search strategies.
@@
This paper is structured as follows: Sec. II reviews simplified leptophilic composite ADM models. Section III examines existing collider constraints. Novel dark shower signatures at $B$ factories and Higgs factories are explored in Secs. IV and V, respectively. Section VI presents our conclusions and future research directions.
@@ II A brief review of QCD-like dark sector with bi-charged mediators
In the leptophilic composite ADM models proposed in Ref. Zhang:2021orr , new heavy particles, including Majorana fermions and $SU(3)_{D}$ charged Dirac fermions, generate out-of-equilibrium processes and explicitly break hidden baryon number conservation. However, these particles lie significantly above current collider energy scales and are consequently decoupled from our analysis. We, therefore, restrict the Lagrangian to the relevant scalar mediator $X$ and dark quark $q_{d}$ interactions as follows [^1] The possible Higgs portal coupling between the SM Higgs doublet field and the $X$ field is irrelevant to this study. Therefore, we assume that its contributions are negligible and can be neglected here.:
@@
Figure 1: Diagrams illustrating dark shower production evolving into invisible or emerging lepton jets. Left panel: $e^{+}e^{-}$ annihilation via $t$-channel mediator $X$ producing a dark quark pair ($q_{d}\overline{q_{d}}$). Right panel: $q\overline{q}$ annihilation yielding mediator pairs, where each mediator decays to a dark quark and a SM lepton.
At lepton and hadron colliders, dark quarks are produced through $t$-channel processes and/or mediator decays, as shown in Fig. 1. Postproduction, they acquire significant kinetic energy ($E\gg\Lambda_{d}$), where $\Lambda_{d}$ denotes the dark confinement scale. Assuming the dark QCD sector mirrors SM QCD dynamics, these dark quarks undergo dark parton shower evolution and hadronization, forming a collinear bunch of dark mesons.
In subsequent sections, we adopt the parameter settings for the QCD-like dark sector from Refs. Schwaller:2015gea ; Mies:2020mzw ; Archer-Smith:2021ntx ; Carrasco:2023loy to enable concrete simulation predictions. We assume three dark colors ($N_{C_{d}}=3$) and seven dark flavors ($N_{f_{d}}=7$), with the following mass relations: $\Lambda_{d}=m_{q_{d}}=2m_{\pi_{d}}$, and $m_{\rho_{d}}=4m_{\pi_{d}}$. Note that the dark confinement scale $\Lambda_{d}$ is closely related to the mass of the lightest dark baryon, which serves as the DM candidate Bai:2013xga ; Schwaller:2015gea ; Renner:2018fhh ; Mies:2020mzw ; Zhang:2021orr . All dark quarks are taken to be mass-degenerate, while dark pions and dark rho mesons are similarly degenerate in both flavor diagonal and off-diagonal states. Other dark baryons and mesons, assumed substantially heavier than dark pions and rho mesons, exhibit negligible production rates from dark showers and hadronization due to mass suppression. We, therefore, exclude them from this analysis.
Integrating out the heavy mediator [final terms in Eq. (1)] yields a dimension-six operator,
@@
which enables dark pion ($\pi_{d}$) decays to charged lepton pairs. The decay width is Schwaller:2015gea
$\Gamma(\pi_{d}\to\bar{\ell}\ell)=\frac{\kappa^{4}f_{\pi_{d}}^{2}m_{\ell}^{2}m_{\pi_{d}}}{32\pi m_{X}^{4}}\sqrt{1-\frac{4m_{\ell}^{2}}{m_{\pi_{d}}^{2}}},\quad\text{with}\quad\ell=e,\mu,\tau.$ (3)
Here $f_{\pi_{d}}$ is the dark pion decay constant and $m_{\ell}$ the lepton mass. Parametrizing the proper lifetime as
$c\tau_{0}=\frac{c\hbar}{\Gamma_{\pi_{d}}}\approx 1.92\,\text{m}\times\left(\frac{1}{\kappa^{4}}\right)\left(\frac{1\,\text{GeV}}{f_{\pi_{d}}}\right)^{2}\left(\frac{0.1\,\text{GeV}}{m_{\ell}}\right)^{2}\left(\frac{1\,\text{GeV}}{m_{\pi_{d}}}\right)\left(\frac{m_{X}}{1\,\text{TeV}}\right)^{4}\left({1-\frac{4m_{\ell}^{2}}{m_{\pi_{d}}^{2}}}\right)^{-\frac{1}{2}}$ (4)
reveals that dark pions propagate measurable distances from the primary vertex before decaying into charged lepton pairs.
@@
To illustrate the typical proper decay length of $\pi_{d}$ versus $m_{\pi_{d}}$, we fix $\kappa=1.0$, $f_{\pi_{d}}=m_{\pi_{d}}$, and $m_{X}=1$ TeV as benchmark values. In order to simplify our discussions here, the lepton flavor universality (LFU) is assumed such that the common coupling $\kappa$ applies to the interactions of mediator with a dark quark and a SM charged lepton for each lepton flavor. Figure 2 shows that for $2m_{e}<m_{\pi_{d}}<2m_{\mu}$, $\pi_{d}$ decays to electron-positron pairs. The proper decay length exceeds $10^{7}$ m, allowing electrons and positrons to escape detection. When $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$, the muon decay channel opens, with $\pi_{d}$ predominantly decaying into muon pairs. The resulting decay length $\sim 1$ m makes these long-lived $\pi_{d}$ detectable. Finally, for $m_{\pi_{d}}>2m_{\tau}$, the tau channel becomes accessible, and $\pi_{d}$ primarily decays to tau pairs, potentially exhibiting prompt decays.
@@ III Existing collider constraints on model parameters
Figure 2 reveals a dramatic change in the proper decay length of $\pi_{d}$ at the $m_{\pi_{d}}\geq 2m_{\mu}$ threshold. We, therefore, classify two dark pion mass ranges: $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$ and $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$, and discuss their respective existing collider constraints.
@@
Figure 3: Exclusion region for dark pion masses $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$. The red dashed line separates long-lived (below) and prompt (above) $X^{\pm}$ decays. The blue and black dashed lines indicate where all $\pi_{d}$ particles escape detection at the LHC and LEP, respectively. The blue and black shaded regions correspond to slepton search constraints (LHC) ATLAS:2019lff and monophoton search constraints (LEP) Fox:2011fx ; Liu:2019ogn . The green region represents the parameter space where the proper lifetime of $\pi_{d}$ exceeds one second, conservatively evaluated at $m_{\pi_{d}}=0.2\text{ GeV}$.
@@
Figure 4: Exclusion region for dark pion masses $2m_{\mu}<m_{\pi_{d}}<2m_{\tau}$. The blue dashed line indicates where all $\pi_{d}$ particles escape detection at the LHC, making this region subject to slepton search constraints. The red dashed line retains the same interpretation as in Fig. 3. The green region represents the parameter space where the proper lifetime of $\pi_{d}$ exceeds one second, conservatively evaluated at $m_{\pi_{d}}=3\text{ GeV}$.
@@
First, we consider the production of scalar mediator pairs through $pp\to X^{+}X^{-}$, with decays $X^{+}\to\ell^{+}q_{d}$ and $X^{-}\to\ell^{-}\overline{q_{d}}$ ($\ell=e,\mu,\tau$). Here, $\kappa$ in Eq. (1) affects only the $X^{\pm}$ lifetime, not the production cross section. When $X^{\pm}$ decays promptly and dark parton showers are invisible at ATLAS and CMS detectors, signatures resemble slepton and stau searches in supersymmetry models ATLAS:2019lff ; ATLAS:2024fub . Under the LFU assumption, we recast these constraints onto this model. We generate a UFO model file for Eq. (1) via FeynRules Alloul:2013bka and simulate signal processes with MadGraph5_aMC@NLO Alwall:2014hca (version 3.5.8). For each $m_{X}$, we scan $\kappa$ and define the mediator prompt decay region by $\gamma\beta c\tau_{X}<0.1\text{ mm}$ in the lab frame. Figures 3 and 4 show regions above the red dashed line $\kappa\gtrsim\mathcal{O}(10^{-7}-10^{-6})$] for the mediator prompt decays recast from slepton and stau constraints [^2] The signal production cross sections rescaled to match Fig. 3 of Ref. Zhang:2021orr with QCD NLO+NLL corrections at $13\text{ TeV}$ LHC Fiaschi:2018xdm .. Additionally, because there are plenty of dark pions produced after dark showers and hadronization, we require the $\pi_{d}$ proper decay length (without any boost) larger than $10\text{ m}$ to conservatively assume all dark pions escape ATLAS and CMS detectors as missing energy. Specifically, we set $m_{\pi_{d}}=f_{\pi_{d}}=0.2\text{ GeV}$ in Fig. 3 and $m_{\pi_{d}}=f_{\pi_{d}}=3\text{ GeV}$ in Fig. 4 to estimate the upper bounds of $\kappa$. Regions below blue dashed lines $\kappa\lesssim\mathcal{O}(1-10)$ for $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$; $\kappa\lesssim\mathcal{O}(0.1-1)$ for $2m_{\mu}\leq m_{\pi_{d}}<2m_{\tau}$] correspond to invisible dark showers which are adapted for recasting.
@@
Under the LFU assumption, the mediator $X^{\pm}$ has approximately equal decay branching ratios of $1/3$ to each flavor channel ($e^{\pm}q_{d}$, $\mu^{\pm}q_{d}$, $\tau^{\pm}q_{d}$). Consequently, we include an extra factor of $1/9$ when applying stau pair production constraints. Comparison with ATLAS results from Table 1 of auxiliary material in Ref. ATLAS:2024fub shows that for $m_{X}>200$ GeV and the dark quark mass $m_{q_{d}}=1$ GeV, the effective cross section is at least a factor of 2 below the experimental bounds. Thus, no exclusion region exists for $m_{X}>200$ GeV, though stau constraints exclude $m_{X}\lesssim 160$ GeV.
@@
For slepton pair production constraints in Fig. 3(c) of Ref. ATLAS:2019lff , the analysis involves all combinations of a pair of leptons from electrons and muons in the final states. Therefore, we apply a factor of $4/9$ in our process. Notably, while the constraints include contributions from both $\tilde{\ell}^{+}_{L}\tilde{\ell}^{-}_{L}$ and $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ production, our signal process specifically corresponds to the $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ channel. We, therefore, isolate the right-handed slepton constraints using the fraction of $pp\to\tilde{\ell}^{+}_{L}\tilde{\ell}^{-}_{L}$ and $pp\to\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ production cross sections from Refs. Fuks:2013lya ; Fiaschi:2018xdm . The resulting effective cross section in our process exceeds ATLAS exclusion limits for $\tilde{\ell}^{+}_{R}\tilde{\ell}^{-}_{R}$ in the mass range $m_{X}\in[200,700]$ GeV. Additionally, after accounting for the average Lorentz boost of the mediator ($\gamma\beta\simeq 1.5$), we estimate the prompt decay boundaries within this mass range.
@@
Conversely, for $\pi_{d}$ with moderate decay lengths, the novel signatures of displaced lepton jets are expected to be observable at the LHC. However, Ref. Zhang:2021orr notes detection challenges for the $pp\to X^{+}X^{-}$ production channel due to (1) insufficient production cross sections, and (2) detection efficiencies after considering the whole cut-flow below $\sim 0.1$. Moreover, significant irreducible cosmic-ray backgrounds further complicate the observation. While LHC Run 3 and HL-LHC may improve sensitivity, current ATLAS displaced lepton jet searches ATLAS:2022izj exclusively target Higgs exotic decays, distinct from this model’s signature of two opposite-sign leptons plus two displaced lepton jets. Consequently, recasting strategies from Ref. ATLAS:2022izj are inapplicable. A dedicated analysis (beyond the scope of this work) would be required, though we establish that the LHC constraints remain weaker than those derived for $e^{+}e^{-}$ colliders in Secs IV and V.